# Addressing Local Realism through Bell Tests at Colliders

### v1

@@ Abstract
One of the most notable aspects of quantum systems is that their components can exhibit correlations much stronger than those allowed by classical physics. Two examples of quantum correlations are quantum entanglement and Bell nonlocality, but generally there is a hierarchy of many types of quantum correlations. Among these correlations, Bell nonlocality holds a special place because it plays a dual role in distinguishing theories where local realism is a valid description. A Bell test, which is a test of local realism, typically needs to be augmented with assumptions to address possible loopholes in the experimental setup. In this work, we study Bell tests in experiments in which the detector reports the correct outcome with a specified probability. This mirrors the situation at high-energy colliders, where particle spins are not measured directly but inferred from the angular distributions of their decay products. We show that, in this setup, a test of local realism is not possible. Quantum correlations, however, are still present, measurable, and informative in high-energy colliders.
@@ I Introduction
Quantum mechanics underpins much of modern physics; however, its nature is still mysterious. The evidence in favor of quantum mechanics is enormous and supported by generations of experiments. Traditional quantum mechanical experiments are performed at low energies in systems of photons or electrons. These systems may consist of two or more particles that exhibit correlations that exceed what is possible from classical physics. A single particle represents a single unit of information, the bit, or for elementary particles whose interactions are governed by quantum mechanics, the qubit.
@@
Recently, collider experiments have begun probing quantum phenomena directly [8]. This started with quantum entanglement between spins in the $t\bar{t}$ system [9], which has already been measured by ATLAS [10] and CMS [11, 12]. Following this, early proposals for measuring Bell nonlocality, which may be measurable in $t\bar{t}$ at the high luminosity run of the Large Hadron Collider (LHC), have been put forward [13, 14, 15, 16]. Other quantum correlations like quantum discord [17], steerability [18], non-negative conditional entropy [17], and the nonlocal advantage of quantum coherence [19] have also been studied in the $t\bar{t}$ final state. A variety of other final states and other high-energy colliders have also been suggested as systems in which to study Bell nonlocality [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 19, 58, 59, 60]. In this work, we exclusively discuss spins as qubits, however, this is an analogous situation when using flavor quantum numbers as qubits [61].
Although early work has investigated the applicability of Bell tests at colliders [62, 63, 64, 65, 66, 67], some confusion persists. In this work we confirm the statement of Refs. [62, 63, 66, 67] that local realism cannot be tested at colliders. We show that local realism cannot be tested with current high-energy particle detectors for two primary reasons. Firstly, there are no independent detector settings, which are required for Bell tests. As a result, Bell’s inequality reduces to differential cross sections, characterized by commuting measurements, which can be described by a local hidden-variable model. Secondly, even if we had independent detector settings, we do not measure the spins of particles directly. The rest-frame decay-product angular distributions can be used to infer the spins of particles, but the rest-frame decay-product angles are not sufficiently strongly correlated with the spin to allow Bell’s inequality to be violated. We demonstrate this with a direct calculation.
@@ II Bell Tests
It is not possible to close every loophole, however, as the number of remaining loopholes shrinks, the class of allowed local hidden variable models becomes increasingly baroque [78]. For that reason, while it is preferable that an experiment address more loopholes, experiments which cannot address some loopholes, like collider experiments, are still interesting.
@@ III.2.2 The General Case
which is a function of $\hat{x}$ measured in the direction $\hat{n}$. Here, $f$ is a function of the scalar quantity $\hat{x}\cdot\hat{n}$.
@@
$\langle\mathcal{O}(\hat{x})_{\hat{n}}\rangle=\frac{1}{2}\Bigl{(}1+f(\hat{x}\cdot\hat{n})\>\hat{n}\cdot\hat{B}\Bigr{)}$, (28)
@@
$+f(\hat{x}_{A}\cdot\hat{\alpha})f(\hat{x}_{B}\cdot\hat{\beta})\>\hat{\alpha}\cdot C\cdot\hat{\beta}\Bigr{)}$.
@@
Unlike the case of an ideal detector, here there is an efficiency factor $\epsilon_{\rm CHSH}=(\Delta F_{A})(\Delta F_{B})$ that dilutes the potential of a state to violate the inequality. [^2] Ref. [83] called the efficiency factor $C$ rather than $\epsilon_{\rm CHSH}$. From Eq. (33) we find
@@
From the previous simple example, we see that the value of $P_{+}^{A}=3/4$ in Eq. (25) leads to the prefactor of $1/4$ in Eq. (24). These yield an efficiency factor of $\epsilon_{\rm CHSH}=1/4$ which means that the CHSH inequality cannot be violated in such a system and consequently that local realism cannot be addressed.
@@ IV.2 Without Assuming Spins
From Eq. (37) we find that $P_{+}=(1/2)(1+\kappa/2)$. We see intuitively that when $\kappa=0$, corresponding to no correlation between the particle spin and the rest-frame decay-product angle, resulting in $\epsilon_{\rm CHSH}\to 0$ meaning that there is no possibility to violate Bell’s inequality. Generally, we find
@@ IV.4 At Future Detectors
It is possible that at future collider experiments the associated detectors will be able to measure the spins of particles [88, 8]. If the spin measurements are direct, as in Sec. III.1, then the experimenter would choose a measurement axis. In this case, Kasday’s local hidden variable model does not apply and local realism can be tested.
@@ IV.5 Beyond Quantum Mechanics
The ideal test would not assume spins and utilize rest-frame decay-product angles directly, as presented in Sec. IV.2. The efficiency factor $\epsilon_{\rm CHSH}$ enters Eq. (43) via $S_{\rm CHSH}\to\epsilon_{\rm CHSH}S_{\rm CHSH}$ which renders the Cirelson bound untestable. Unlike with local realism, while testing the Cirelson bound it is not invalidating to assume the underlying description is built on spins. This case, however, is also untestable because of the measurement method applied at colliders.
@@
One possibility to indicate superquantum behavior would be deviations in the distributions used to extract the Fano coefficients. Distributions that do not match the quantum field theory predictions, if verified, could point to a superquantum theory.
@@ V Entanglement and Bell Nonlocality
As noted in Refs. [62, 63, 66, 67], local realism cannot be tested in collider experiments because only cross sections are measured, allowing a local hidden variable model to account for all observed data. [^5] Refs. [62, 63, 66, 67] also state that the reason cross sections have a local hidden variable model is that the measured final state momenta commute. It has been shown that commuting measurements can still lead to tests of local realism if the choice of detector settings was made at some point in the past and transmitted, along with the results, to the experimenters [91]. In Sec. III and Sec. IV we have demonstrated this in an alternative way, namely that due to the distribution of decay products, the efficiency factor multiplying the CHSH inequality is too low to test local realism.
@@ VI Conclusions
This was captured by an efficiency factor that dilutes the observation of the terms in the CHSH inequality relative to the threshold required for violation. We showed that for a Bell test to be possible the efficiency factor must be greater than $1/\sqrt{2}$. At high-energy colliders, the largest possible factor is $1/4$, rendering this an ineffective test. Future detectors that could measure particle spins directly would be able to test local realism because experiments in which spin is directly measured have an efficiency factor of $1$.

### v2

@@ Abstract
One of the most notable aspects of quantum systems is that their components can exhibit correlations much stronger than those allowed by classical physics. Two examples of quantum correlations are quantum entanglement and Bell nonlocality, but generally there is a hierarchy of many types of quantum correlations. Among these correlations, Bell nonlocality holds a special place because it plays a dual role in distinguishing theories where local realism is a valid description. A Bell test, which is a test of local realism, typically needs to be augmented with assumptions to address possible loopholes in the experimental setup. In this work, we study Bell tests in experiments in which the detector reports the correct outcome with a specified probability. This mirrors the situation at high-energy colliders, where particle spins are not measured directly but inferred from the angular distributions of their decay products. We show that, in this setup, a test of local realism is not possible. Quantum correlations, however, are still present, measurable, and informative in high-energy colliders. These correlations are the building blocks of the interesting, developing quantum information science program at high-energy colliders. The measurements of entanglement by the ATLAS and CMS experiments are the first steps in this initiative.
@@ I Introduction
Quantum mechanics underpins much of modern physics; however, its nature is still mysterious. The evidence in favor of quantum mechanics is enormous and supported by generations of experiments [1, 2, 3, 4, 5, 6]. Traditional quantum mechanical experiments are performed at low energies in systems of photons or electrons. These systems may consist of two or more particles that exhibit correlations that exceed what is possible from classical physics. A single particle represents a single unit of information, the bit, or for elementary particles whose interactions are governed by quantum mechanics, the qubit [7].
@@
Recently, collider experiments have begun probing quantum phenomena directly [15]. This started with quantum entanglement between spins in the $t\bar{t}$ system [16], which has already been measured by ATLAS [17] and CMS [18, 19]. Following this, early proposals for measuring Bell nonlocality, which may be measurable in $t\bar{t}$ at the high luminosity run of the Large Hadron Collider (LHC), have been put forward [20, 21, 22, 23]. Other quantum correlations like quantum discord [24], steerability [25], non-negative conditional entropy [24], and the nonlocal advantage of quantum coherence [26] have also been studied in the $t\bar{t}$ final state. A variety of other final states and other high-energy colliders have also been suggested as systems in which to study Bell nonlocality [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 26, 65, 66, 67, 68]. In this work, we exclusively discuss spins as qubits, however, this is an analogous situation when using flavor quantum numbers as qubits. In this case, the different decay times neutral meson systems allow different correlations to be probed. Unfortunately, local realism cannot be tested using flavor in current experiments [69, 70].
Although early work has investigated the applicability of Bell tests at colliders [71, 72, 73, 74, 75, 76], some confusion persists. In this work we confirm the statement of Refs. [71, 72, 75, 76] that local realism cannot be tested at colliders. We show that local realism cannot be tested with current high-energy particle detectors for two primary reasons. Firstly, there are no independent detector settings, which are required for Bell tests. As a result, Bell’s inequality reduces to differential cross sections, characterized by commuting measurements, which can be described by a local hidden-variable model. Secondly, even if we had independent detector settings, we do not measure the spins of particles directly. The rest-frame decay-product angular distributions can be used to infer the polarizations and spin correlations of particles, but the inherent distribution of these angles is too wide to be used as a proxy for the spins. Consequently, while strong spin correlations may exist in the underlying quantum state, they are diluted when comparing the correlations between rest-frame decay angles. As a result, Bell’s inequality cannot be violated in this system. We demonstrate this with a direct calculation.
@@ II Bell Tests
It is not possible to close every loophole, however, as the number of remaining loopholes shrinks, the class of allowed local hidden variable models becomes increasingly baroque [83]. Superdeterminism is a loophole that always persists. For that reason, while it is preferable that an experiment address more loopholes, experiments which cannot address some loopholes, like collider experiments, are still interesting.
@@ III.2.2 The General Case
which is a function of $\hat{x}$ measured in the direction $\hat{n}$. Here, $f$ is an odd function of the scalar quantity $\hat{x}\cdot\hat{n}$.
@@
[…]
@@
[…]
@@
Unlike the case of an ideal detector, here there is a theoretical efficiency factor $\epsilon_{\rm CHSH}=(\Delta F_{A})(\Delta F_{B})$ that dilutes the potential of a state to violate the inequality. [^2] Ref. [88] called the theoretical efficiency factor $C$ rather than $\epsilon_{\rm CHSH}$. This is not an experimental efficiency that can be corrected, but rather a theoretical efficiency that results from the measurement operator. From Eq. (33) we find
@@
From the previous simple example, we see that the value of $P_{+}^{A}=3/4$ in Eq. (25) leads to the prefactor of $1/4$ in Eq. (24). These yield a theoretical efficiency factor of $\epsilon_{\rm CHSH}=1/4$ which means that the CHSH inequality cannot be violated in such a system and consequently that local realism cannot be addressed.
@@ IV.2 Without Assuming Spins
From Eq. (37) we find that $P_{+}=(1/2)(1+\kappa/2)$. We see intuitively that when $\kappa=0$, corresponding to no spin information transfer between the particle spin and the rest-frame decay-product angle, resulting in $\epsilon_{\rm CHSH}\to 0$ meaning that there is no possibility to violate Bell’s inequality. Generally, we find
@@ IV.4 At Future Detectors
It is possible that at future collider experiments the associated detectors will be able to measure the spins of particles [93, 15]. If the spin measurements are direct, as in Sec. III.1, then the experimenter would choose a measurement axis. In this case, Kasday’s local hidden variable model does not apply and local realism can be tested, when the measurement axes can be changed as needed.
@@ IV.5 Beyond Quantum Mechanics
The ideal test would not assume spins and utilize rest-frame decay-product angles directly, as presented in Sec. IV.2. The theoretical efficiency factor $\epsilon_{\rm CHSH}$ enters Eq. (43) via $S_{\rm CHSH}\to\epsilon_{\rm CHSH}S_{\rm CHSH}$ which renders the Cirelson bound untestable. Unlike with local realism, while testing the Cirelson bound it is not invalidating to assume the underlying description is built on spins. This case, however, is also untestable because of the measurement method applied at colliders.
@@
One possibility to indicate superquantum behavior would be deviations in the distributions used to extract the Fano coefficients. Distributions that do not match the quantum field theory predictions, if verified, could point to a superquantum theory (see, for example, Ref. [96]).
@@ V Entanglement and Bell Nonlocality
As noted in Refs. [71, 72, 75, 76], local realism cannot be tested in collider experiments because only cross sections are measured, allowing a local hidden variable model to account for all observed data. [^5] Refs. [71, 72, 75, 76] also state that the reason cross sections have a local hidden variable model is that the measured final state momenta commute. It has been shown that commuting measurements can still lead to tests of local realism if the choice of detector settings was made at some point in the past and transmitted, along with the results, to the experimenters [97]. In Sec. III and Sec. IV we have demonstrated this in an alternative way, namely that due to the distribution of decay products, the theoretical efficiency factor multiplying the CHSH inequality is too low to test local realism.
@@ VI Conclusions
This was captured by a theoretical efficiency factor that dilutes the observation of the terms in the CHSH inequality relative to the threshold required for violation. We showed that for a Bell test to be possible the theoretical efficiency factor must be greater than $1/\sqrt{2}$. At high-energy colliders, the largest possible factor is $1/4$, rendering this an ineffective test. Future detectors that could measure particle spins directly would be able to test local realism because experiments in which spin is directly measured have a theoretical efficiency factor of $1$.